OpenStudy (anonymous):

I WILL GIVE MEDAL AND FAN!!! Variable y varies directly with variable x, and y = 20 when x = 8. Enter the constant of variation in lowest terms. y = __/__ x

OpenStudy (igreen):

Direct variation: y = kx Plug in what we know: 20 = k(8) 20 = 8k Divide 8 to both sides

OpenStudy (anonymous):

is it 5/2?

OpenStudy (igreen):

Yep, you got it. 20/8 = 5/2

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (igreen):

2.5, 2 1/2

OpenStudy (igreen):

No problem, Happy to Help \(\Huge\ddot\smile\)